Suppose the utility function for Bernard is given by Z=2H1H2 and his income equation by Y=P1H1+P2H2

Derive the demand equation for H1H2

1
Expert's answer
2021-05-17T10:41:58-0400

given utility function "Z=2H_1H_2"

income constraint "Y=P_1H_1+P_2H_2"

utility maximization condition "MRS=\frac{P_1}{P_2}"

"MRS=\frac {\delta H_1}{\delta H_2}=\frac{2H_2}{2H_2}=\frac{H_2}{H_2}"

"\frac{H_2}{H_1}=\frac{P_1}{P_2}"

"H_2-\frac{P_1}{P_2}H_1......(I)"

putting equation (i) in the income constraint

"Y=P_1H_1+P_2(\frac{P_1}{P_2}H_1)"

"Y=P_1H_1+P1H_1"

"Y=2P_1H_1"

"H_1=\frac{Y}{2P_1}" .....(Demand equation for "H_1" )


Putting "H_1" in equation (I)

"H_2=\frac{P_1}{P_2}(\frac{Y}{2P_1})"

"H_2=\frac{2}{2p_2}" ......(demand function for "H_2" )